40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5243     Entry time: Mon Aug 15 21:43:29 2011
Author: Anamaria and Keiko 
Type: Summary 
Category: Locking 
Subject: central part ifo locking project 

 REFL33 and REFL165 cables were connected from the AP table to the rack.  Cables on the rack for REFL33I, 33Q, 165I, 165Q ports were connected, too. Connections were confirmed by the data viewer. Two SMA cables which will be used for the two PDs on the AP tabl were built. We will be able to place the two PDs tomorrow. The beamsplitters to split the laser to REFL33 and REFL165 ports were mounted and ready to be placed.

ELOG V3.1.3-